OpenStudy (anonymous):

solve the diff equation. (y^-1)dy + (y*e^cos(x))*(sin(x))dx

OpenStudy (anonymous):

Equals zero?

OpenStudy (anonymous):

yes ... =0

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

It looks separable to me.

OpenStudy (anonymous):

i got it to \[- e ^{\cos(x)} = y^{-1} +C\] is this correct first off?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i cannot figure out how to solve for y

OpenStudy (anonymous):

Put the arbitrary constant on the other side. It is arbitrary, so it doesn't matter which side of the equation it is on.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

but it is y^-1

OpenStudy (anonymous):

Well, if we have\[y^{-1}=C-e^{\cos(x)}\]We can simply take the reciprocal of both sides, provided that it doesn't equal zero:\[y=\frac{1}{C-e^{\cos(x)}}\]

OpenStudy (anonymous):

alright thanks that's the answer.. couldn't figure out how to change it but it makes sense